OpenStudy (anonymous):

-3x/4 - 4= x/6 how do i solve?

OpenStudy (anonymous):

-(3 x)/4-4 = x/6 -(11 x)/12 = 4 x = -48/11

OpenStudy (anonymous):

-3/4(x)-1/6(x) -9/12 - 2/12 = -11/12 -11x/12 =4 -11x = 48 x=-48/11